OpenStudy (anonymous):

In anna's pocket there is $16.20 in quarters, nickles and dimes. In all there are 97 coins if there are seven more nickles than dimes, find how many of each kind there are.

OpenStudy (anonymous):

25 quaarter 97 - 2x+7 pcs 5 = nickle x+7 pcs 10 = dimes x pcs 1620 total 97pcs 25(104-2x) + 5(x+ 7) + 10x = 1620 2600 - 50x + 5x + 35 + 10x = 1620 -35x = 1620 - 2635 -35x = -1015 x= 29 x+ 7 = 36 29+ 36 = 65 97- 65 = 32 there are 29 dimes 36 nickles and 32quarters